Find the mid points of the line joining the pairs of points
a.(3,4) and (5,2)
b.(0,6) and (4,0)
c.(4,2) and (4,-4)
d.(-1,-5) and (-3,4)
e.(-3,2) and (8,-2)
1
Expert's answer
2020-07-16T19:07:30-0400

a. "(\frac{3+5}{2},\frac{4+2}{2})=(4,3)"

b. "(\frac{0+4}{2},\frac{6+0}{2})=(2,3)"

c. "(\frac{4+4}{2},\frac{2-4}{2})=(4,-1)"

d. "(\frac{-1-3}{2},\frac{-5+4}{2})=(-2,-\frac{1}{2})"

e. "(\frac{-3+8}{2},\frac{2-2}{2})=(\frac{5}{2},0)"